# The TAP equation: evaluating combinatorial innovation in biocosmology

### v3

@@ 1 Introduction
The Theory of the Adjacent Possible (TAP) [1] is the notion that the near-future outcomes of some developmental process are limited by the objects that already exist. The objects under study might be physical (molecules, gene sequences, species, etc.) or conceptual (patents, memes, songs, etc.) The TAP equation provides a simple model for the adjacent possible by counting the ways that new objects can be generated from combinations of existing objects. It allows for an efficiency in creation, so that not all possible new objects are realised at each stage. It is assumed that at each timestep, new objects can be created from any combination of the existing objects, while some fraction of existing objects expire (extinction or obsolescence). It is therefore a model of combinatorial innovation [2].
@@
Here the $\alpha_{i}$ are a set of decreasing constants accounting for the increasing difficulty in linking up larger numbers of elements, the final term is the combinatorial combinations of existing elements, and $10\leq\mu<$ 1 is the extinction rate of existing objects. As it stands Eq. (1) is not quite well-defined due to $M_{t}$ not being constrained to be an integer, but this can be fixed either by analytic interpolation (for instance replacing factorials by Gamma functions) or stochasticity [4].
The equation permits dramatically-explosive behaviour, much faster than exponential, because not only do the combinatoric terms rapidly become vast, but so too does the number of terms in the summation. Indeed typical solutions grow so rapidly as to effectively diverge to infinity in a finite time. Technically they do not diverge at a finite point in time as every term in the summation is finite, despite the number of combinations soon becoming unimaginably large. But this can be considered an artefact of the discretisation and we will see that any analytic approximation to the solution diverges in a finite time, as seen in Ref. [4]. We refer to this divergence as ‘blow-up’.
The TAP equation is under active investigation as a possible model for a variety of processes. Koppl et al. [5] study three different phenomena: the explosive recent growth of GDP per capita, the rapidly-developing diversity of manufactured goods, and the family tree of US patent applications. In Ref. [6] we consider TAP as a model for an evolving state space in the Universe due to emergent laws associated with the appearance of life (see also Ref. [7]). Generally, it may be applicable to the collection of phenomena describing the climate indicators — loosely termed ‘The Great Acceleration’ [8]. It can also be said to model the the ongoing impacts of civilisation on our environment which were first termed ‘Technological Singularity’ in 1993 by NASA’s conference proceedings article in Ref. [9] (part of NASA’s legacy Cloud-based Data Match-up Service (CDMS) collection). The same phenomenon is also described in the more recent work in Ref. [10].
In this article we obtain a series of new mathematical results for several variants of the TAP equation, that may be relevant across all its applications.
2 Solving the TAP equation
@@
First we make a couple of technical observations. For our results in this section we start the sum in Eq. (1) at $i=2$, not $i=1$ as in Steel et al. [4], in order that the pre-existing elements are not duplicated at each step. However, in an alternative view including the $i=1$ term allows for new items to be created through evolution from a single starting point, and we study this and other TAP variants in Section 4. More significantly, we note that Steel et al.’s stochastic implementation is not a discretisation of the TAP equation itself, because it requires a calculational timestep to be small to keep the creation probabilities below unity; it is a stochastic discretization of a continuum approximation to the original equation. Their numerical analysis also makes the further assumption of a fixed upper limit (usually $i_{\rm max}=4$) in the summation. As we deal directly with the discrete TAP equation, we see some detailed differences in results.
@@
When the $\alpha_{i}$ are equal, $\alpha_{i}=\alpha$, the sum in Eq. (1) can be carried out analytically; it is just $\alpha$ times the sum of the corresponding row of Pascal’s triangle minus its first two entries, which is $12^{M_{t}}-M_{t}-$ 1. The TAP equation now becomes
$M_{t+1}=M_{t}(1-\mu)+\alpha\left(2^{M_{t}}-M_{t}-1\right)$ (2)
whose behaviour can more or less be read off by comparing the relative sizes of the two terms. Except for special values of $\mu$ and $\alpha$ this formula will generate non-integer values but it can be taken as an analytic continuation to those values (the continuation allowing both for the combinatoric elements and the range of the summation being ill-defined for non-integer $M_{t}$).
To see the extraordinarily rapid growth this equation implies, we first set $\alpha$ equal to 1 and the extinction $\mu$ to zero. Starting with $M_{0}=2$, the sequence goes like this:
@@
Using $\alpha=1$ and $\mu=0$ means we are counting all the possible combinations that can arise via the adjacent possible. In later implementations only a subset are generated.
Only five steps has generated a number much larger than the number of particles in the observable Universe! As the number is shifted into the exponent at each step, a specialist notation for extremely large numbers known as tetration, described in Section 3, is required to provide an analytic expression.
[…]
$t_{\rm double}=\frac{M_{0}}{\alpha\left(2^{M_{0}}-M_{0}-1\right)}$ (3)
@@
$t_{\rm add\;one}=\frac{1}{\alpha\left(2^{M_{0}}-M_{0}-1\right)}\,.$ (4)
@@
Figure 1: Critical values of $\mu$ as a function of the parameters $M_{0}$ and $\alpha$. From top to bottom the contours are $\mu_{\rm critical}$ equal to $1$, $10^{-2}$, $10^{-4}$, $10^{-6}$. The shaded region is where extinction equilibrium is impossible as it would require $\mu>1$.
@@
$2$ $1$ 00 5 2
$2$ $10^{-2}$ 00 85 200
$2$ $10^{-4}$ 00 7800 20000
$2$ $10^{-2}$ $10^{-3}$ 94 200
$2$ $10^{-2}$ $10^{-2}$ Never Never
$10$ $10^{-2}$ 00 4 0.2
$10$ $10^{-4}$ 00 20 20
$2$ $10^{-2}$ $4\times 10^{-3}$ 167
$2$ $10^{-2}$ $4.9\times 10^{-3}$ 327
$2$ $10^{-2}$ $4.99\times 10^{-3}$ 504
$2$ $10^{-2}$ $4.999\times 10^{-3}$ 686
@@
such that each additional element to be combined reduces the success rate by a factor of $a$ [^1] Steel et al.’s simulations in Ref. [4] use $a=100$, but also cut the sum off at $i_{\rm max}=4$. At least until the vicinity of blow-up when discretisation effects become important, the $\alpha$ parameter is essentially degenerate with the units of the timestepping, provided the extinction is also suitably scaled. For instance if we choose e.g. $\alpha=0.1$ we simply get ten times as many outputs from the code in reaching the same ultimate outcome. It would hence be natural to now set it to one, but for completeness we retain $\alpha$ for now.
@@
$t_{\rm double}$ $=$ $\frac{M_{0}}{\alpha a\left[\left(1+1/a\right)^{M_{0}}-M_{0}/a-1\right]}\,;$ (9)
@@
$2$ $1$ 00 5 1
$2$ $10$ 00 18 10
$2$ $100$ 00 144 100
$2$ $1000$ 00 1392 1000
$6$ $10000$ 00 3600 2000
$10$ $10$ 00 5 1
$10$ $100$ 00 25 11
$10$ $1000$ 00 216 111
$100$ $10000$ 00 206 100
Table 2: Blow-up times for the varying $\alpha_{i}$ model, with $\alpha=1$. The analytic estimate is Eq. (12), whose proportionalities with $a$ and $M_{0}$ are clearly seen in the numerical results.
@@ 4.1 The ‘two-scale’ TAP equation
where $\alpha_{1}$ is the rate of evolution from individual objects and $\alpha\ll\alpha_{1}$ is the multi-object combination rate from before, now highly suppressed relative to the evolution term. We see immediately that the evolution rate is perfectly degenerate with the extinction $\mu$, hence all results go through as before, except that we are now motivated to consider $\alpha_{1}>\mu$ that corresponds to the previously-impossible case of anti-extinction.
@@ 4.2 The differential TAP equation
The TAP equation with $\alpha=1$ and no extinction seeks to count all adjacent possibilities. However the equation regenerates the same items over and over at each step, so that for instance an object first created at step 2 is then duplicated at each subsequent step from the same constituents. We can correct for this simply by subtracting off the number of items created at the previous step. Taking care to retain the original object set $M_{0}$ the equation now reads
@@ 4.3 The logistic TAP equation
We mention one further idea which we reserve for future investigation: the question of whether there is a logistic TAP equation. In population dynamics, the logistic mapping arises when the birth rate is moderated by making the parents compete for resources, most simply modelled by adding a $-M^{2}$ term [12]. This model already shows period doubling and beautiful deterministic chaos.

### v4

@@ 1 Introduction
The Theory of the Adjacent Possible (TAP) [1] is the notion that the near-future outcomes of some developmental process are limited by the objects that already exist. The objects under study might be physical (molecules, gene sequences, species, etc.) or conceptual (patents, memes, songs, etc.) The TAP equation provides a simple model for the adjacent possible by counting the ways that new objects can be generated from combinations of existing objects. It allows for an efficiency in creation, so that not all possible new objects are realised at each stage. It is assumed that at each timestep, new objects can be created from any combination of existing objects, while some fraction of existing objects expire (extinction or obsolescence). It is therefore a model of combinatorial innovation [2].
@@
Here the $\alpha_{i}$ are a set of decreasing constants accounting for the increasing difficulty in linking up larger numbers of elements, the final term is the combinatorial combinations of existing elements, and $0\leq\mu\leq 1$ is the extinction rate of existing objects. As it stands Eq. (1) is not quite well-defined due to $M_{t}$ not being constrained to be an integer, but this can be fixed either by analytic interpolation (for instance replacing factorials by Gamma functions) or stochasticity [4].
The equation permits dramatically-explosive behaviour, much faster than exponential, because not only do the combinatoric terms rapidly become vast, but so too does the number of terms in the summation. Indeed, typical solutions grow so rapidly as to effectively diverge to infinity in a finite time. Technically they do not diverge at a finite point in time as every term in the summation is finite, despite the number of combinations soon becoming unimaginably large. But this can be considered an artefact of the discretisation and we will see that any analytic approximation to the solution diverges in a finite time, as seen in Ref. [4]. We refer to this divergence as ‘blow-up’.
The TAP equation is under active investigation as a possible model for a variety of processes. Koppl et al. [5] study three different phenomena: the explosive recent growth of GDP per capita, the rapidly-developing diversity of manufactured goods, and the family tree of US patent applications. In Ref. [6] we consider TAP as a model for an evolving state space in the Universe due to emergent laws associated with the appearance of life (see also Ref. [7]). Generally, it may be applicable to the collection of phenomena describing the climate indicators — loosely termed ‘The Great Acceleration’ [8]. It can also be said to model the ongoing impacts of rapid technological development which were first termed ‘Technological Singularity’ in 1993 by NASA’s conference proceedings article in Ref. [9] (part of NASA’s legacy Cloud-based Data Match-up Service (CDMS) collection). The same phenomenon is also described in the more recent work in Ref. [10].
In this article we obtain a series of new mathematical results for several variants of the TAP equation, which may be relevant across all its applications.
2 Approximating solutions to the TAP equation
@@
First we make a couple of technical observations. For our results in this section we start the sum in Eq. (1) at $i=2$, not $i=1$ as in Steel et al. [4], in order that the pre-existing elements are not duplicated at each step. However, in an alternative view including the $i=1$ term allows for new items to be created through evolution from a single starting point, and we study this and other TAP variants in Section 5.
More significantly, we note that Steel et al.’s stochastic implementation is not a discretisation of the TAP equation itself, because it requires a calculational timestep to be small to keep the creation probabilities below unity; it is a stochastic discretization of a continuum approximation to the original equation. Moreover, non-linearity means that even the mean stochastic behaviour will depart from the deterministic evolution. Their numerical analysis also makes the further assumption of a fixed upper limit (usually $i_{\rm max}=4$) in the summation. As we deal directly with the discrete TAP equation, we see some detailed differences in results.
@@
When the $\alpha_{i}$ are equal, $\alpha_{i}=\alpha$, the sum in Eq. (1) can be carried out analytically; it is just $\alpha$ times the sum of the corresponding row of Pascal’s triangle minus its first two entries, which is $2^{M_{t}}-M_{t}-1$. The TAP equation now becomes
$M_{t+1}=M_{t}(1-\mu)+\alpha\left(2^{M_{t}}-M_{t}-1\right)\,$, (2)
whose behaviour can more or less be read off by comparing the relative sizes of the two terms. Except for special values of $\mu$ and $\alpha$ this formula will generate non-integer values of $M_{t}$ but it can be taken as an analytic continuation to those values (the continuation allowing both for the combinatoric elements and the range of the summation being ill-defined for non-integer $M_{t}$).
To see the extraordinarily rapid growth this equation implies, we first set $\alpha$ equal to 1 and the extinction $\mu$ to zero. Starting with two elements, $M_{0}=2$, the sequence goes as follows:
@@
Using $\alpha=1$ and $\mu=0$ means we are counting all the possible combinations that can arise via the adjacent possible. In later implementations, only a subset are generated.
Only five steps have generated a number much larger than the number of particles in the observable Universe! As the number is shifted into the exponent at each step, a specialist notation for extremely large numbers, known as tetration, described in Section 3, is required to provide an analytic expression.
Blow-up happens as soon as the second term dominates the first, so that each step dramatically changes the total, which then massively enhances the next addition. This is typically when $M_{t}\,2^{-M_{t}}<\alpha$. If this is already satisfied by $M_{0}$ at $t=0$ blow-up is immediate, otherwise there is initially a plateau of slow growth. As a first estimate of the length of the plateau we might consider the doubling-time as estimated from the initial growth rate:
$t_{\rm double}\simeq\frac{M_{0}}{\alpha\left(2^{M_{0}}-M_{0}-1\right)}$ (3)
@@
$t_{\rm add\;one}\simeq\frac{1}{\alpha\left(2^{M_{0}}-M_{0}-1\right)}\,.$ (4)
@@
Figure 1: Critical values of $\mu$ as a function of the parameters $M_{0}$ and $\alpha$. From top to bottom the contours are $\mu_{\rm critical}$ equal to 11, $10^{-2}$, $10^{-4}$, $10^{-6}$. The shaded region is where extinction equilibrium is impossible as it would require $\mu>1$.
@@
22 11 0 5 2
22 $10^{-2}$ 0 85 200
22 $10^{-4}$ 0 7800 20000
22 $10^{-2}$ $10^{-3}$ 94 200
22 $10^{-2}$ $10^{-2}$ Never Never
$10$ $10^{-2}$ 0 4 0.2
$10$ $10^{-4}$ 0 20 20
22 $10^{-2}$ $4\times 10^{-3}$ 167
22 $10^{-2}$ $4.9\times 10^{-3}$ 327
22 $10^{-2}$ $4.99\times 10^{-3}$ 504
22 $10^{-2}$ $4.999\times 10^{-3}$ 686
@@ 22 $10^{-2}$ $4\times 10^{-3}$ 167
such that each additional element to be combined reduces the success rate by a factor of aa. [^1] Steel et al.’s simulations in Ref. [4] use $a=100$, but also cut the sum off at $i_{\rm max}=4$. At least until the vicinity of blow-up when discretisation effects become important, the $\alpha$ parameter is essentially degenerate with the units of the timestepping, provided the extinction is also suitably scaled. For instance if we choose e.g. $\alpha=0.1$ we simply get ten times as many outputs from the code in reaching the same ultimate outcome. It would hence be natural to now set it to one, but for completeness we retain $\alpha$ for now.
@@
$t_{\rm double}$ $\simeq$ $\frac{M_{0}}{\alpha a\left[\left(1+1/a\right)^{M_{0}}-M_{0}/a-1\right]}\,;$ (9)
@@
22 11 0 5 1
22 $10$ 0 18 10
22 $100$ 0 144 100
22 $1000$ 0 1392 1000
66 $10000$ 0 3600 2000
$10$ $10$ 0 5 1
$10$ $100$ 0 25 11
$10$ $1000$ 0 216 111
$100$ $10000$ 0 206 100
Table 2: Blow-up times for the varying $\alpha_{i}$ model, with $\alpha=1$. The analytic estimate is Eq. (12), whose proportionalities with $a$ and $M_{0}$ are clearly seen in the numerical results (defined as in Table 1).
@@ 3 Late-time TAP
4 More on blow-up
We now make some further remarks on blow-up. As already explained, the situation is rather different depending on whether one views the discrete TAP equation as fundamental or as an approximation to a continuous process. In the former case, blow-up is not completely well-defined, as even despite the late-time tetration the evolution does not diverge in a finite number of steps. However, it does not seem reasonable to approximate any process as discrete on a given timescale once it begins evolving rapidly on that timescale. We proposed an operational definition of the time to blow-up using the timescale to add one item and then adding a series of such steps, e.g. equation (12) for the power-law $\alpha_{i}$ case, which gives an upper limit on when the discrete model will become unjustifiable.
In cases where an analytic continuation of the $M_{t+1}$ formula to real-valued $M(t)$ exists, such as the power-law $\alpha_{i}$, this can be made more rigorous by taking the continuous limit $dt\rightarrow 0$. [^2] We thank one of the referees, Mike Steel, for providing this insight. We then write
$\frac{dM(t)}{dt}=F[M]\,$, (15)
where $F[M]$ is the analytically-continued function that gives the evolution, including any specified model parameters such as the $\alpha_{i}$ [e.g. the right-hand side of equation (8)]. The time to reach infinity from an initial value $M_{0}$ is then given by the integral
$t_{{\rm blow-up}}=\int_{M_{0}}^{\infty}\frac{dM}{F[M]}\,.$ (16)
For a sufficiently-rapidly growing function $F[M]$, which TAP always gives, this will yield a finite value. This provides a rigorously-defined estimate of the effective blow-up time of the discrete model, which can also be interpreted as a time before which the discrete timestepping model must lose viability due to the rate of evolution.
As an example, this formula applied to the fourth row of Table 2 given a blow-up time estimate of 1381, very close to the computed discrete-model blow-up time.
Further investigations of the blow-up are given in this volume by Bellina et al. [12], who connect the phenomenon to a number of real-world systems.
@@ 5.1 The ‘two-scale’ TAP equation
where $\alpha_{1}$ is the rate of evolution from individual objects and $\alpha\ll\alpha_{1}$ is the multi-object combination rate from before, now highly suppressed relative to the evolution term. We see immediately that the evolution rate is perfectly degenerate with the extinction $\mu$, hence all results go through as before, except that we are now motivated to consider $\alpha_{1}>\mu$ that corresponds to the previously impossible case of anti-extinction.
@@ 5.2 The differential TAP equation
The TAP equation with $\alpha=1$ and no extinction seeks to count all adjacent possibilities. However, the equation regenerates the same items over and over at each step, so that, for instance, an object first created at step 2 is then duplicated at each subsequent step from the same constituents. We can correct for this simply by subtracting the number of items created at the previous step. Taking care to retain the original object set $M_{0}$ the equation now reads
@@ 5.3 The logistic TAP equation
We mention one further idea which we reserve for future investigation: the question of whether there is a logistic TAP equation. In population dynamics, the logistic mapping arises when the birth rate is moderated by making the parents compete for resources, most simply modelled by adding a $M^{2}$ term [13]. This model already shows period doubling and beautiful deterministic chaos.